OpenStudy (firejay5):

An apple in a tree has a gravitational potential energy of 175 J and a weight of 3.5 N. How high from the ground is the apple?

OpenStudy (firejay5):

Who is right? Teacher's Answers: 4. 50 m My Answers 4. h = PE/mg = 175/(3.5)(9.8) = 5.1m

OpenStudy (anonymous):

You mistakenly used the weight of the apple (3.5 N) as the mass. This error is a very common one. The mg in your solution is 3.5 N.

OpenStudy (firejay5):

I need clarification on how to show my work for this problem

OpenStudy (anonymous):

Well, we know that \[F_g = mg = 3.5 \text{ N}\]Therefore, \[\Delta h = \frac{ E_g }{ mg }=\frac{ E_g }{ F_g }\]\[=\frac{ 175 }{ 3.5 }\]\[=5.0 \times 10^1 \text{ m}\]

OpenStudy (firejay5):

okay nevermind @ospreytriple thanks for your help on all of my Physics work it means a lot to me

OpenStudy (anonymous):

You're welcome

OpenStudy (firejay5):

no one rarely helps me on here with the best explanation